# On functoriality and the tensor product property in noncommutative tensor-triangular geometry

### v1

@@ Abstract.
Noncommutative, tensor-triangular geometry, support variety, complete prime, triangulated category, Balmer spectrum, tensor product property, functorial
@@ 1. Introduction
But what about the oft-occurring monoidal triangulated categories for which the tensor product is not necessarily symmetric? Analogously to the differences between standard algebraic geometry vs. noncommutative algebraic geometry (see [Rey25]), things are less clear-cut. Support varieties for settings without a symmetric tensor product (or any tensor product at all) have been studied for quite some time [EHS+03, FW11, Lin10, SS04] and Buan, Krause, and Solberg laid some of the framework for developing noncommutative tensor-triangular geometry via theories of support using frame theory in [BKS07, BKSS20]. However, only recently a noncommutative analogue of the Balmer spectrum was formalized by Nakano, Vashaw, and Yakimov in [NVY22a], and further examined by the authors in [NVY22b, NVY23, NVY24]. A similar construction was adopted by Negron and Pevtsova in [NP23].
@@
More generally, we give a criteria for precisely when a monoidal exact functor $F:\mathscr{K}\to\mathscr{L}$ induces a well-defined map on Balmer spectra.
Theorem (Corollary 4.3).
Let $F\colon\mathscr{K}\to\mathscr{L}$ be an exact $\otimes$-functor, and denote its image by $F(\mathscr{K})\subseteq\mathscr{L}$. Then $F$ induces a continuous map on Balmer spectra $\varphi:=\operatorname{Spc}(F):\operatorname{Spc}(\mathscr{L})\to\operatorname{Spc}(\mathscr{K})$ given by $\mathscr{Q}\mapsto F^{-1}(\mathscr{Q})$ if and only if for all prime thick $\otimes$-ideals $\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})$, the intersection $F(\mathscr{K})\cap\mathscr{Q}$ is a prime thick $\otimes$-ideal of $F(\mathscr{K})$.
This is perhaps discouraging, but there is more to the picture. First, the authors propose alternative axioms of support to consider, encouraging the notion that there may not be one fixed notion of support data, akin to how there are multiple spectral theories for noncommutative rings, or how support axioms break down for big tensor-triangulated categories. In their companion piece [NVY22b], the authors analyze when $(\operatorname{Spc}(\mathscr{K}),\operatorname{supp})$ satisfies the tensor product property, i.e. $\operatorname{supp}(a)\cap\operatorname{supp}(b)=\operatorname{supp}(a\otimes b)$. They prove this property holds if and only if every prime is completely prime, i.e. a prime $\mathscr{P}$ which satisfies the “commutative prime” condition, $a\otimes b\in\mathscr{P}$ implies $a\in\mathscr{P}$ or $b\in\mathscr{P}$. The question of whether the the tensor product property holds is for certain cohomological support varieties is of independent interest and has been examined further in [NP23, BPW24, BPW25].
If we think of this complete prime spectrum $\operatorname{Spc}^{cp}(\mathscr{K})$ as another construction of a spectral space or support theory, then in some ways it is better-behaved; for instance, it is indeed functorial. However, it has its own major shortcomings (as explained in [NVY22a]) - notably, it may be empty, and the standard “prime lifting” techniques (see [Bal05, Lemma 2.2] or [NVY22a, Theorem 3.2.3]) do not seem feasible if $\operatorname{Spc}^{cp}(\mathscr{K})\subset\operatorname{Spc}(\mathscr{K})$. However, in tensor-triangular geometry, one should think about the Balmer spectrum in the terms of universal support: the (commutative) Balmer spectrum is the universal support datum for tensor-triangulated categories, and the noncommutative Balmer spectrum is universal for, in fact, multiple notions of noncommutative support data. And similarly, the complete prime spectrum is itself universal - it is the final object in the category of multiplicative support data, i.e. support data satisfying the tensor product property.
@@
Let $\mathscr{S}^{cp}(\mathscr{K})$ denote the category of multiplicative support data (see Definition 2.7 and 3.1). The following hold.
  • (Theorem 3.2) $\mathscr{S}^{cp}(-):\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Cat}$ is a contravariant functor from the category of essentially small monoidal triangulated categories to the category of essentially small categories.
  • (Theorem 4.5) The complete prime spectrum $(\operatorname{Spc}^{cp}(\mathscr{K}),\operatorname{supp}^{cp})$ is the final object of $\mathscr{S}^{cp}(\mathscr{K})$, even if $\operatorname{Spc}^{cp}(\mathscr{K})=\emptyset$.
@@
Given a spectral theory on a monoidal triangulated category $\mathscr{K}$, the specialization-closed (or Thomason) subsets should be expected to parametrize some data about $\mathscr{K}$. We show that with some modifications, $\operatorname{Spc}^{cp}(\mathscr{K})$ parametrizes certain thick tensor-ideals.
@@
Finally, following the blueprint laid out by Reyes [Rey12], we expand our scope and determine the universal functorial support theory for monoidal triangulated categories that coincides with $\operatorname{Spc}$ for tensor-triangulated categories, the partial prime spectrum. Moreover, the partial prime spectrum universal in the sense of support axioms - it is the final commutative support data.
@@
The paper is organized as follows. Section 2 discusses preliminaries in noncommutative tensor-triangular geometry and proves a few topological properties analogous to the symmetric setting. Section 3 introduces the category of support data and proves (partial) functoriality, and contains discussion about multiplicative support axioms. Section 4 proves functoriality and universality of the complete prime spectrum and partial functoriality of the noncommutative Balmer spectrum. Section 5 proves an analogue to Balmer’s classification of thick tensor-ideals for complete primes. Section 6 is a brief aside: we deduce the final object in the category of quasi-support data, as defined in [NVY22a].
In Section 7, we apply the support theories and Balmer spectra we have considered to the crossed product category $\mathscr{K}\rtimes G$, an extension of Benson and Witherspoon’s crossed product algebra [BW14]. In the last two sections, we switch to analyzing the functoriality of the complete prime spectrum, with the additional assumption that one of our categories is duo, i.e. every one-sided ideal is two-sided. In Section 8, we introduce this notion, discuss rigidity, and discuss the work of Negron-Pevtsova [NP23]. In Section 9, we adapt some of Balmer’s surjectivity results [Bal18] to the noncommutative setting for the complete prime spectrum. Finally in Section 10, we define the partial prime spectrum and determine that it is the universal functorial support theory for monoidal triangulated categories, analogous to Reyes [Rey12].
Acknowledgments: The author thanks Julia Pevtsova for numerous conversations which led to certain key aspects of the paper, Kent Vashaw and Dan Nakano for discussion on aspects of noncommutative tensor-triangular geometry, and Beren Sanders for his detailed feedback on an earlier version of the paper. He also is indebted to Paul Balmer and Martin Gallauer, whose correspondence and conversations throughout the past year helped him quickly enter the world of tensor-triangular geometry.
@@ 2. Preliminaries
We begin by reviewing standard noncommutative tensor-triangular preliminaries, noncommutative analogues of tensor-triangular geometry developed in [Bal05]. We follow the conventions of [NVY22b]. Unless otherwise stated, we assume $\mathscr{K}$ and $\mathscr{L}$ are essentially small monoidal triangulated categories.
@@ 3. Categories of support data
The category $\mathscr{S}(\mathscr{K})$ has a full subcategory $\mathscr{S}^{cp}(\mathscr{K})$, the category of multiplicative support data on $\mathscr{K}$.
By [NVY22a, Theorem 4.2.2], the final object of $\mathscr{S}(\mathscr{K})$ is the Balmer spectrum of $\mathscr{K}$, $\operatorname{Spc}(\mathscr{K})$. Allowing $(\emptyset,\operatorname{id})$ to be a valid support datum endows $\mathscr{S}(\mathscr{K})$ with an initial object. We use the notation $\mathscr{S}(-)$ to suggest functoriality. For $\mathscr{S}^{cp}(-)$, this is indeed the case.
@@
We have a contravariant functor $\mathscr{S}^{cp}(-)\colon\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Cat}$ from the category of essentially small monoidal triangulated categories to the category of essentially small categories. It is defined as follows:
  • An essentially small monoidal triangulated category $\mathscr{K}$ is mapped to its category of multiplicative support data $\mathscr{S}^{cp}(\mathscr{K})$;
  • An exact $\otimes$-functor $F\colon\mathscr{K}\to\mathscr{L}$ is mapped to the following covariant pullback functor $F^{*}\colon\mathscr{S}^{cp}(\mathscr{L})\to\mathscr{S}^{cp}(\mathscr{K})$:
@@
We must verify that $F^{*}(f)$ is a morphism of support data, that is, for all $x\in\mathscr{K}$, $F^{*}(f)^{-1}(\tau\circ F)(x)=(\sigma\circ F)(x)$, however this follows easily via the construction. Finally, it is clear that $F^{*}(\operatorname{id})=\operatorname{id}$ and for $f\colon(Y,\sigma)\to(Z,\tau)$ and $g\colon(Z,\tau)\to(Z^{\prime},\tau^{\prime})$ morphisms of support data, that $F^{*}(g)\circ F^{*}(f)=F^{*}(g\circ f)$, hence $F^{*}$ is functorial, and thus the functor $\mathscr{S}^{cp}(-)$ is well-defined. Functoriality of $\mathscr{S}^{cp}(-)$ amounts to showing $\operatorname{id}^{*}=\operatorname{id}$ and $(G\circ F)^{*}=F^{*}\circ G^{*}$ which is straightforward. ∎
@@
  • An essentially surjective exact $\otimes$-functor $F\colon\mathscr{K}\to\mathscr{L}$ is mapped to the following pullback functor $F^{*}\colon\mathscr{S}^{cp}(\mathscr{L})\to\mathscr{S}^{cp}(\mathscr{K})$:
@@
One must wonder if there is a all-encompassing notion of support. In the ring-theoretic case, Reyes [Rey12] determines that, if one wants the spectrum to be functorial and coincide with commutative ring spectra for commutative rings, one cannot guarantee the spectrum will always be nonempty. Similarly, we want that our support axioms should always coincide with the usual (multiplicative) support axioms for symmetric or braided monoidal categories. We ask the following question: does a universal, functorial, necessarily non-empty support theory exist? Section 10 lays out the groundwork to solve this answer in the same manner as Reyes, but at present the answer remains unclear. [GS23] also hints that the answer is no, though in a much more general setting.
@@
We have presented here a general approach to evaluating different axioms of support, but for the context of this paper we reserve $\mathscr{S}(\mathscr{K})$ and $\mathscr{S}^{cp}(\mathscr{K})$ for the support axioms defined in Definition 3.1, and mainly focus on understanding multiplicative or functorial support data. Throughout this article, we will aim to answer the above questions.
4. Functoriality and universality of the complete prime spectrum
@@
$=\varphi^{-1}(\operatorname{supp}_{\mathscr{K}}(x))$
As a result, we can actually determine exactly when an exact $\otimes$-functor $F:\mathscr{K}\to\mathscr{L}$ induces a well-defined map on noncommutative Balmer spectra. Note that the image of $F$ is a subcategory of $\mathscr{L}$, although not necessarily full, thick, or replete. Note that for the intersection of two subcategories $\mathscr{I}\cap\mathscr{P}$ of $\mathscr{K}$, the intersection is considered on both objects and morphisms, i.e. a morphism $f\in\mathscr{K}$ belongs to $\mathscr{I}\cap\mathscr{P}$ if and only if $f\in\mathscr{I}$ and $f\in\mathscr{P}$. Therefore, the intersection of a full and thick subcategory of $\mathscr{K}$ with any other subcategory is again full and thick.
@@
Let $F\colon\mathscr{K}\to\mathscr{L}$ be an exact $\otimes$-functor, and denote its image by $F(\mathscr{K})\subseteq\mathscr{L}$. Then $F$ induces a continuous map on Balmer spectra $\varphi:=\operatorname{Spc}(F):\operatorname{Spc}(\mathscr{L})\to\operatorname{Spc}(\mathscr{K}),\mathscr{Q}\mapsto F^{-1}(\mathscr{Q})$ if and only if for all prime thick $\otimes$-ideals $\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})$, the intersection $F(\mathscr{K})\cap\mathscr{Q}$ is a prime thick $\otimes$-ideal of $F(\mathscr{K})$.

### v2

@@ Abstract.
Noncommutative, tensor-triangular geometry, support variety, complete prime, triangulated category, Balmer spectrum, tensor product property
@@ 1. Introduction
But what about the oft-occurring monoidal triangulated categories for which the tensor product is not necessarily symmetric? Analogously to the differences between standard algebraic geometry vs. noncommutative algebraic geometry (see [Rey25] for a reader-friendly overview), things are less clear-cut. Support varieties for settings without a symmetric tensor product (or any tensor product at all) have been studied for quite some time [EHS+03, FW11, GS23, Lin10, SS04] and Buan, Krause, and Solberg laid some of the framework for developing noncommutative tensor-triangular geometry via theories of support using frame theory in [BKS07, BKSS20]. However, only recently a noncommutative analogue of the Balmer spectrum was formalized by Nakano, Vashaw, and Yakimov in [NVY22a], and further examined by the authors in [NVY22b, NVY23, NVY24]. A similar construction was adopted by Negron and Pevtsova in [NP23].
@@
More generally, we give a criteria for precisely when a monoidal exact functor $F:\mathscr{K}\to\mathscr{L}$ induces a well-defined map on Balmer spectra. As a result, we can deduce the noncommutative Balmer spectra of localizations and idempotent completion, adaptations of standard tensor-triangular geometric results for the noncommutative setting.
Theorem (Theorem 4.4).
Let $F\colon\mathscr{K}\to\mathscr{L}$ be an exact $\otimes$-functor, and denote its image by $F(\mathscr{K})\subseteq\mathscr{L}$. Then $F$ induces a continuous map on Balmer spectra $\varphi:=\operatorname{Spc}(F):\operatorname{Spc}(\mathscr{L})\to\operatorname{Spc}(\mathscr{K})$ given by $\mathscr{Q}\mapsto F^{-1}(\mathscr{Q})$ if and only if for all prime thick $\otimes$-ideals $\mathscr{Q}\in\operatorname{Spc}(\mathscr{L})$, the intersection $F(\mathscr{K})\cap\mathscr{Q}$ satisfies the primality property,
$\text{For all $a,b\in F(K)$, }a\otimes F(K)\otimes c\in F(K)\cap\mathscr{Q}\text{ implies }a\in\mathscr{Q}\text{ or }b\in\mathscr{Q}.$
The lack of functoriality is perhaps discouraging, but there is more to the picture. First, Nakano, Vashaw, and Yakimov propose alternative axioms of support to consider, encouraging the notion that there may not be one fixed notion of support data, akin to how there are multiple spectral theories for noncommutative rings, or how support axioms break down for big tensor-triangulated categories. In their companion piece [NVY22b], the authors analyze when $(\operatorname{Spc}(\mathscr{K}),\operatorname{supp})$ satisfies the tensor product property, i.e. $\operatorname{supp}(a)\cap\operatorname{supp}(b)=\operatorname{supp}(a\otimes b)$. They prove this property holds if and only if every prime is completely prime, i.e. a prime $\mathscr{P}$ which satisfies $a\otimes b\in\mathscr{P}$ implies $a\in\mathscr{P}$ or $b\in\mathscr{P}$. The question of whether the the tensor product property holds for cohomological support varieties is of independent interest and has been examined further in [NP23, BPW24, BPW25].
If we think of this complete prime spectrum $\operatorname{Spc}^{cp}(\mathscr{K})$ as another construction of a spectral space or support theory, then in some ways it is better-behaved; for instance, it is indeed functorial. However, it has its own major shortcomings (as explained in [NVY22a]) - notably, it may be empty, and the standard “prime lifting” techniques (see [Bal05, Lemma 2.2] or [NVY22a, Theorem 3.2.3]) do not seem feasible if $\operatorname{Spc}^{cp}(\mathscr{K})\subset\operatorname{Spc}(\mathscr{K})$. Frame-theoretically, it is insufficient. However, in tensor-triangular geometry, one should think about the Balmer spectrum in the terms of universal support: the (commutative) Balmer spectrum is the universal support datum for tensor-triangulated categories, and the noncommutative Balmer spectrum is universal for, in fact, multiple notions of noncommutative support data. And similarly, the complete prime spectrum is itself universal - it is the final object in the category of multiplicative support data, i.e. support data satisfying the tensor product property.
@@
Let $\mathscr{S}^{\otimes}(\mathscr{K})$ denote the category of multiplicative support data (see Definition 2.7 and 3.1). The following hold.
  • (Theorem 3.2) $\mathscr{S}^{\otimes}(-):\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Cat}$ is a contravariant functor from the category of essentially small monoidal triangulated categories to the category of essentially small categories.
  • (Theorem 5.2) The complete prime spectrum $(\operatorname{Spc}^{cp}(\mathscr{K}),\operatorname{supp}^{cp})$ is the final object of $\mathscr{S}^{\otimes}(\mathscr{K})$, even if $\operatorname{Spc}^{cp}(\mathscr{K})=\emptyset$.
@@
Given a spectral theory on a monoidal triangulated category $\mathscr{K}$, the specialization-closed or Thomason subsets should be expected to parametrize some data about $\mathscr{K}$. We show that with some modifications, $\operatorname{Spc}^{cp}(\mathscr{K})$ parametrizes certain thick tensor-ideals.
@@
Finally, following the blueprint laid out by Reyes [Rey12], we expand our scope and determine the universal functorial support theory for monoidal triangulated categories that coincides with $\operatorname{Spc}$ for tensor-triangulated categories, the partial prime spectrum. Moreover, the partial prime spectrum universal in the sense of support axioms - it is the final commutative support datum (see Definition 11.14).
@@
The paper is organized as follows. Section 2 discusses preliminaries in noncommutative tensor-triangular geometry and proves a few topological properties analogous to the symmetric setting. Section 3 introduces the category of support data, proves its partial functoriality, and contains discussion about multiplicative support axioms. Section 4 gives a criterion for when monoidal exact functors induce maps on noncommutative Balmer spectrum, and deduces a few standard yet important properties about the spectrum with regards to localization and idempotent completion. Section 5 proves functoriality and universality of the complete prime spectrum. Section 6 proves an analogue to Balmer’s classification of thick tensor-ideals for complete primes. Section 7 is a brief aside: we deduce the final object in the category of quasi-support data, as defined in [NVY22a].
In Section 8, we apply the support theories and Balmer spectra we have considered to the crossed product category $\mathscr{K}\rtimes G$, an extension of Benson and Witherspoon’s crossed product algebra [BW14]. In the next two sections, we switch to analyzing the functoriality of the complete prime spectrum, with the additional assumption that one of our categories is duo, i.e. every one-sided ideal is two-sided. In Section 9, we introduce this notion, discuss rigidity, and discuss the work of Negron-Pevtsova [NP23]. In Section 10, we adapt some of Balmer’s surjectivity results [Bal18] to the noncommutative setting for the complete prime spectrum. Finally in Section 11, we define the partial prime spectrum and determine that it is the universal functorial support theory for monoidal triangulated categories, analogous to Reyes [Rey12].
Acknowledgments: The author thanks Julia Pevtsova for numerous conversations which led to certain key aspects of the paper, Kent Vashaw, Dan Nakano, and Timothy de Deyn for discussions on aspects of noncommutative tensor-triangular geometry, and Beren Sanders for his detailed feedback on an earlier version of this paper. He also is indebted to Paul Balmer and Martin Gallauer, whose correspondence and conversations throughout the past year helped him quickly enter the world of tensor-triangular geometry.
@@ 2. Preliminaries
We begin by reviewing standard noncommutative tensor-triangular preliminaries, noncommutative analogues of tensor-triangular geometry developed in [Bal05]. We follow the conventions of [NVY22b].
@@ 3. Categories of support data
The category $\mathscr{S}(\mathscr{K})$ has a full subcategory $\mathscr{S}^{\otimes}(\mathscr{K})$, the category of multiplicative support data on $\mathscr{K}$.
By [NVY22a, Theorem 4.2.2], the final object of $\mathscr{S}(\mathscr{K})$ is the Balmer spectrum of $\mathscr{K}$, $\operatorname{Spc}(\mathscr{K})$. Allowing $(\emptyset,\operatorname{id})$ to be a valid support datum endows $\mathscr{S}(\mathscr{K})$ with an initial object. We use the notation $\mathscr{S}(-)$ to suggest functoriality. For $\mathscr{S}^{\otimes}(-)$, this is indeed the case.
@@
We have a contravariant functor $\mathscr{S}^{\otimes}(-)\colon\mathbf{mon}_{\Delta}^{\operatorname{op}}\to\mathbf{Cat}$ from the category of essentially small monoidal triangulated categories to the category of essentially small categories. It is defined as follows:
  • An essentially small monoidal triangulated category $\mathscr{K}$ is mapped to its category of multiplicative support data $\mathscr{S}^{\otimes}(\mathscr{K})$;
  • An exact $\otimes$-functor $F\colon\mathscr{K}\to\mathscr{L}$ is mapped to the following covariant pullback functor $F^{*}\colon\mathscr{S}^{\otimes}(\mathscr{L})\to\mathscr{S}^{\otimes}(\mathscr{K})$:
@@
We must verify that $F^{*}(f)$ is a morphism of support data, that is, for all $x\in\mathscr{K}$, $F^{*}(f)^{-1}(\tau\circ F)(x)=(\sigma\circ F)(x)$, however this follows easily via the construction. Finally, it is clear that $F^{*}(\operatorname{id})=\operatorname{id}$ and for $f\colon(Y,\sigma)\to(Z,\tau)$ and $g\colon(Z,\tau)\to(Z^{\prime},\tau^{\prime})$ morphisms of support data, that $F^{*}(g)\circ F^{*}(f)=F^{*}(g\circ f)$, hence $F^{*}$ is functorial, and thus the functor $\mathscr{S}^{\otimes}(-)$ is well-defined. Functoriality of $\mathscr{S}^{\otimes}(-)$ amounts to showing $\operatorname{id}^{*}=\operatorname{id}$ and $(G\circ F)^{*}=F^{*}\circ G^{*}$ which is straightforward. ∎
@@
  • An essentially surjective exact $\otimes$-functor $F\colon\mathscr{K}\to\mathscr{L}$ is mapped to the following pullback functor $F^{*}\colon\mathscr{S}^{\otimes}(\mathscr{L})\to\mathscr{S}^{\otimes}(\mathscr{K})$:
@@
One must wonder if there is a all-encompassing notion of support. In the ring-theoretic case, Reyes [Rey12] determines that, if one wants the spectrum to be functorial and coincide with commutative ring spectra for commutative rings, one cannot guarantee the spectrum will always be nonempty. Similarly, we want that our support axioms should always coincide with the usual (multiplicative) support axioms for symmetric or braided monoidal categories. We ask the following question: does a universal, functorial, necessarily non-empty support theory exist? The final section lays out the groundwork to solve this answer in the same manner as Reyes, but at present the answer remains unclear.
@@
We have presented here a general approach to evaluating different axioms of support, but for the context of this paper we reserve $\mathscr{S}(\mathscr{K})$ and $\mathscr{S}^{\otimes}(\mathscr{K})$ for the support axioms defined in Definition 3.1, and mainly focus on understanding multiplicative or functorial support data. Throughout this article, we will aim to answer the above questions.
4. Partial functoriality of the noncommutative Balmer spectrum
@@
$=\varphi^{-1}(\operatorname{supp}_{\mathscr{K}}(x)).$
Given any thick $\otimes$-ideal $\mathscr{I}$ of $\mathscr{K}$, one can perform the localization at $\mathscr{I}$, i.e. taking the Verdier quotient $\mathscr{K}/\mathscr{I}$, which is again monoidal triangulated. We have a canonical localization functor $\mathscr{K}\twoheadrightarrow\mathscr{K}/\mathscr{I}$, which is surjective on objects.
@@
Let $\mathscr{I}$ be a thick $\otimes$-ideal of $\mathscr{K}$, and let $q\colon\mathscr{K}\to\mathscr{K}/\mathscr{I}$ be a localization. The map $\varphi:=\operatorname{Spc}(q)\colon\operatorname{Spc}(\mathscr{K}/\mathscr{I})\to\operatorname{Spc}(\mathscr{K})$ induces a homeomorphism between $\operatorname{Spc}(\mathscr{K}/\mathscr{I})$ and the subspace $\{\mathscr{P}\in\operatorname{Spc}(\mathscr{K})\mid\mathscr{I}\subseteq\mathscr{P}\}$ of $\operatorname{Spc}(\mathscr{K})$ of prime thick $\otimes$-ideals containing $\mathscr{I}$. Similarly, the map $\varphi^{cp}:\operatorname{Spc}^{cp}(q)\colon\operatorname{Spc}^{cp}(\mathscr{K}/\mathscr{I})\to\operatorname{Spc}(\mathscr{K})$ induces a homeomorphism between $\operatorname{Spc}(\mathscr{K}/\mathscr{I})$ and the subspace $\{\mathscr{P}\in\operatorname{Spc}^{cp}(\mathscr{K})\mid\mathscr{I}\subseteq\mathscr{P}\}$ of $\operatorname{Spc}^{cp}(\mathscr{K})$.